An ideal gas in the amount of 0, 9 moles first expands isobarrically and then cools isochorically. During the entire process, an amount of heat 2000 J is transferred to the gas. The initial and final gas temperatures are the same and equal to 260 K. Find the ratio of the final gas volume to its initial volume.
1
Expert's answer
2020-11-27T00:40:57-0500

After the isochoric cooling, the temperature is 260 K, the heat transferred to the gas is 0 (no work done and it is cooling). During the process of isobaric expansion the heat transferred is


"Q=P\Delta V=PV_2-PV_1.\\
PV_1=nRT_1,\\
PV_2=nRT_2.\\\space\\
Q=nRT_2-nRT_1,\\\space\\
T_2=\frac{Q}{nR}+T_1."


From the first process of isobaric expansion, we have


"\frac{V_2}{V_1}=\frac{T_2}{T_1},\\\space\\
\frac{V_2}{V_1}=\frac{Q}{nRT_1}+1=2.03."